# Unifying Multitrack Music Arrangement via Reconstruction Fine-Tuning and Efficient Tokenization

### v2

@@ 6 Conclusion
In this paper, we proposed a unified framework for automatic music arrangement that addresses challenges in tokenization, pre-trained model utilization, and arrangement quality. Our tokenizer demonstrates superior efficiency and enhances modeling effectiveness across various generation tasks, while our training objective shows compatibility with diverse music arrangement tasks. Further, LM pre-training significantly enhances downstream performance, particularly for tasks requiring complex sequence alignment. Taken together, our method achieves state-of-the-art results on band arrangement, piano reduction, and drum arrangement tasks, generating natural, faithful, coherent, creative, and musically appealing outputs. The performance advantage of our framework suggests broader applications in music-to-music transformation tasks like infilling and call-and-response generation. Furthermore, our tokenizer’s compact and efficient design makes it a standalone contribution applicable to general multitrack symbolic music generation tasks beyond the scope of arrangement.
@@ A.1 Vocabulary
Table 1 presents the complete vocabulary of our proposed REMI-z tokenizer, detailing the value ranges for each token type. For instrument tokens, the values correspond to MIDI program numbers, with 128 specifically designated for the drum set. For pitched instruments, pitch token values directly map to MIDI pitch numbers, while drum set pitches are encoded as MIDI pitch + 128. Both position and duration tokens are quantized in units of a 48th note (one-third of a sixteenth note). Time signature and tempo tokens undergo specific quantization before token mapping; for detailed quantization rules, please refer to our code implementation. All tokens listed in the table were utilized during pre-training, with example REMI-z sequences illustrated in Figure 3(b).
During fine-tuning, we simplified the token set by excluding time signature and tempo tokens. This decision was supported by our analysis of the fine-tuning dataset (Slakh2100 Manilow et al. [2019]), where 94.79% of songs use 4/4 time signatures. Consequently, we restricted our fine-tuning to 4/4 songs and omitted time signature tokens, following practices in related works Wang et al. [2022]; Zhao et al. [2024]. We also excluded tempo tokens since tempo adjustments in digital audio workstations are typically handled as a global parameter, affecting only tempo tokens without altering the any other tokens. This simplification assumes that compositional and performance styles remain consistent across different tempos. However, this assumption may not hold for datasets with significant tempo variations. Therefore, for future research requiring time-signature- or tempo-specific characteristics, we recommend including these tokens during fine-tuning.
Velocity tokens were excluded from both pre-training and fine-tuning phases. This decision reflects our focus on compositional quality rather than performance naturalism, aligning with previous approaches in band and piano arrangement studies Zhao et al. [2024]; Terao et al. [2023]. However, if velocity information is deemed crucial for generation, our model architecture readily accommodates the addition of velocity tokens to each note without introducing content fragmentation issues, if the REMI-z note organization order is maintained.
@@ A.2 Example
(a) A 4-bar musical segment with orange arrows illustrating the ‘zig-zag’ encoding order of notes in REMI-z tokenization.
@@
We name our tokenization scheme REMI-z for its distinctive “zig-zag” encoding pattern for musical notes within each bar, as illustrated in Figure 3(a). In this scheme, notes are encoded hierarchically: first grouped by tracks (instruments), then organized bar by bar. This track-first approach ensures notes from the same instrument remain clustered together, thereby enhancing the model’s ability to learn instrument-specific patterns. The resulting REMI-z sequence is demonstrated in Figure 3(b). In contrast, REMI+ von Rütte et al. [2023] employs a column-wise encoding strategy, strictly ordering notes by their temporal positions. While this approach effectively captures global temporal relationships, it disperses notes from the same instrument throughout the sequence, potentially complicating instrument-specific pattern learning. We leverage this temporal ordering advantage of REMI+ in our content sequence (excluding instrument tokens), as shown in Figure 4.
@@ B.1 Key Normalization
Transposition-equivariance is a crucial property in symbolic music: a composition’s musicality remains unchanged under global pitch shifts (uniform pitch adjustment across all notes). However, uneven key distribution in datasets can lead to data sparsity during training, potentially causing models to perform inconsistently across different keys and incorrectly assume key-dependent composition rules. Two approaches address this issue: (1) data augmentation through systematic semitone transpositions ${0,\pm 1,\pm 2,\cdots}$ Malandro [2024]; Zhao et al. [2024], or (2) normalizing all songs to a common key (e.g., C major and A minor) Lu et al. [2023]. We adopt the latter approach, implementing a modified version of Lu et al. [2023]’s method. Our implementation uses a key dictionary mapping 24 keys (12 major and 12 minor) to 12-dimensional binary vectors, where 1s indicate scale notes. Key detection is performed by computing dot products between these vectors and a song’s pitch histogram, with the highest-scoring key determining the transposition needed to normalize to C major or A minor. Further details can be found in the code.
B.2 Instrument Quantization
@@ B.3 Baseline Models
When comparing the effectiveness of REMI+ and the proposed REMI-z tokenization, the model structure used is the same as our arrangement model, but with a simpler training objective—the standard left-to-right language modeling.
@@
For band arrangement, we use the Transformer-VAE model from Zhao et al. [2024] as our baseline, utilizing their official implementation without retraining. The model’s generation module was trained on Slakh2100 Manilow et al. [2019] (the same dataset as our fine-tuning) and its prior model leveraged the larger Lakh MIDI Dataset Raffel [2016], which encompasses Slakh2100.
@@
Piano reductions can be categorized into two types: (1) piano accompaniments where the melody is delegated to a separate lead instrument, focusing solely on preserving harmony and texture (e.g., Zhao et al. [2023]; Wang et al. [2020]), which can be effectively handled using self-supervised methods; and (2) complete solo piano arrangements that additionally include the melody, requiring careful human arrangement and band-to-piano parallel datasets for supervised learning (e.g., Terao et al. [2023, 2022]). Due to the lack of open-source parallel datasets, we focus on the accompaniment arrangement task in this paper and use the term piano reduction interchangeably. However, our methodology is inherently flexible and can also handle the second type of reduction if parallel datasets become available. Additionally, exploring the potential for full solo piano reductions using unsupervised approaches remains a intriguing direction for future research.
For comparison, we reimplemented the UNet baseline from Terao et al. [2023] and trained it on Slakh2100 using our data preparation pipeline. Notably, we omitted octave shifts during input processing for two reasons: First, our output consistently represents a subset of the input notes, unlike their setup where human-composed piano references might use different octaves for the same notes. Second, introducing octave shifts in training significantly degraded performance, reducing note F1 from 58.31 to 42.80 and diminishing perceptual quality.
@@
For drum arrangement, we employ Composer’s Assistant v2 Malandro [2024] as our baseline, specifically using v2.1.0 from their official repository [^2] https://github.com/m-malandro/composers-assistant-REAPER. We use the model without retraining since its original training data (Lakh MIDI Dataset Raffel [2016]) encompasses our fine-tuning dataset (Slakh2100).
B.4 Hyperparameter Settings
@@
For fine-tuning, we conducted a simple learning rate search over 1e-5, 5e-5, 1e-4, selecting the optimal rate based on validation loss. This resulted in learning rates of 5e-5 for drum arrangement and 1e-4 for band arrangement and piano reduction. The batch sizes and context lengths were configured as follows: band arrangement used a batch size of 24 and context length of 768; piano reduction similarly adopted a batch size of 24 and context length of 768; drum arrangement employed a batch size of 8 and context length of 1536. Across all fine-tuning tasks, we used the AdamW optimizer with 0.01 weight decay, incorporating a linear learning rate scheduler with 500-step warmup. Training spanned 3 epochs for band, piano, and drum tasks, with early stopping patience of 2 epochs. The best checkpoints were selected based on validation loss performance.
@@ Appendix C Objective Metrics
In this section, we formally define the objective metrics used in the paper.
Most of our objective metrics are based on note-level F1 scores calculated on piano roll of 16-th note quantization. Let $X=\{x_{1},...,x_{n}\}$ and $Y=\{y_{1},...,y_{m}\}$ be two sequences of note events, where each note event $e$ is defined as a tuple $e=(t,p)$ with onset time $t$ and pitch $p$. A note event $x_{i}\in X$ is considered to match $y_{j}\in Y$ if and only if:
@@
[…]
@@
[…]
where $S$, $D$, and $I$ are the minimum number of substitutions, deletions, and insertions required to transform $V^{out}$ into $V^{tgt}$, and $N=|V^{tgt}|$ is the length of the target sequence.
@@ D.2 Human Evaluation Details
For the band arrangement task, we use an out-of-domain test with new compositions and new instrument groups for subjective evaluation, while for piano reduction and drum arrangement, we use songs from the test set to facilitate comparison with the original piano track (Rule-O) and the ground truth drum track. Each model/method was tasked with arranging the full song, and a phrase from the chorus, the most representative part, was selected for comparison, each containing 15s–30s of audio. We prepared 6 songs for the band arrangement, 5 for the piano reduction, and 5 for the drum arrangement. In the questionnaire, participants were first presented with the original song clip to be arranged, followed by the anonymized results from different models/methods in a random order. Participants were asked to evaluate each arranged clip on a 5-point scale, using the questions provided for each metric as outlined in §Subjective Metrics. The group of songs evaluated by each participant (containing one original arrangement and all corresponding model outputs) was also randomly selected. In total, we received 56, 73, and 77 groups of feedback for the band arrangement, piano reduction, and drum arrangement tasks, respectively.
Here are specific details for the band arrangement setting. In the band arrangement tasks, we tested the models’ ability in different scenarios by using various instrument settings. The content conditions were derived from either a piano arrangement or a full band arrangement. The instrument conditions included three different instrument sets, arranged in descending pitch order: string trio (violin, viola, cello), rock band (synth lead, clean electric guitar, distorted electric guitar, electric bass), and jazz band (saxophone, violin, brass section, clean electric guitar, piano, string ensemble, electric bass). Among the 6 songs used for the band arrangement task, two songs were arranged in each of the three settings. Additionally, the input came from songs composed for different settings, including 3 piano solos and 3 band arrangements.

### v3

@@ 6 Conclusion
We presented a unified framework for automatic music arrangement, centered around a reconstruction objective that enables diverse arrangement tasks without requiring parallel data, enabling knowledge transfer from generative pre-trained symbolic music models for arrangement tasks. Our approach also integrates a structured tokenization scheme, REMI-z, which convert multitrack music to compact and easy-to-model token sequences. Experimental results on band arrangement, piano reduction, and drum arrangement show that our method consistently outperforms task-specific baselines in both objective and subjective evaluations. These results demonstrate the potential of our framework as a general solution for symbolic music-to-music transformation.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
@@ A.1 Vocabulary
Table 6: REMI-z vocabulary with pitch token distinctions
Meaning Token X’s range
[…]
End of a bar b-1 -
[…]
Table 6 presents the complete vocabulary of our proposed REMI-z tokenizer, detailing the value ranges for each token type. For instrument tokens, the values correspond to MIDI program numbers (0–127), with 128 specifically designated for the drum set. For pitched instruments, pitch token values directly map to MIDI pitch numbers, while drum set pitches are encoded as MIDI pitch + 128. Both position and duration tokens are quantized in units of a 48th note (one-third of a sixteenth note). Time signature and tempo tokens undergo specific quantization before token mapping; for detailed quantization rules, please refer to our code implementation. All tokens listed in the table were utilized during pre-training, including time signature and tempo tokens serving as bar-level properties, with example REMI-z sequences illustrated in Figure 4(b).
During fine-tuning, we simplified the token set by excluding time signature and tempo tokens. This decision was supported by our analysis of the fine-tuning dataset (Slakh2100 [19]), where 94.8% of songs use 4/4 time signatures. Consequently, we restricted our fine-tuning to 4/4 songs and omitted time signature tokens, following practices in related works [32, 39]. We also excluded tempo tokens since tempo adjustments in digital audio workstations are typically handled as a global parameter, affecting only tempo tokens without altering the any other tokens. This simplification assumes that compositional and performance styles remain consistent across different tempos. However, this assumption may not hold for datasets with significant tempo variations. Therefore, for future research requiring time-signature- or tempo-specific characteristics, we recommend including these tokens during fine-tuning.
Velocity tokens were excluded from both pre-training and fine-tuning phases. This decision reflects our focus on compositional quality rather than performance naturalism, aligning with previous approaches in band and piano arrangement studies [39, 27]. However, if velocity information is deemed crucial for generation, our model architecture readily accommodates the addition of velocity tokens to each note without introducing content fragmentation issues, if the REMI-z note organization order is maintained.
@@ A.2 Example
(a) A 4-bar musical segment with orange arrows illustrating the “zig-zag” encoding order of notes in REMI-z tokenization.
@@
We name our tokenization scheme REMI-z for its distinctive “zig-zag” encoding pattern for musical notes within each bar, as illustrated in Figure 4(a). In this scheme, notes are encoded hierarchically: first grouped by tracks (instruments), then organized bar by bar. This track-first approach ensures notes from the same instrument remain clustered together, thereby enhancing the model’s ability to learn instrument-specific patterns. The resulting REMI-z sequence is demonstrated in Figure 4(b). In contrast, REMI+ [28] employs a column-wise encoding strategy, strictly ordering notes by their temporal positions. While this approach effectively captures global temporal relationships, it disperses notes from the same instrument throughout the sequence, potentially complicating instrument-specific pattern learning. We leverage this temporal ordering advantage of REMI+ in our content sequence (excluding instrument tokens), as shown in Figure 5.
@@ B.1 Key Normalization
Transposition-equivariance is a crucial property in symbolic music: a composition’s musicality remains unchanged under global pitch shifts (uniform pitch adjustment across all notes). However, uneven key distribution in datasets can lead to data sparsity during training, potentially causing models to perform inconsistently across different keys. Two approaches address this issue: (1) data augmentation through systematic semitone transpositions $0,\pm 1,\pm 2,\cdots$} [17, 39], or (2) normalizing all songs to a common key (e.g., C major and A minor) [15]. We adopt the latter approach, implementing a modified version of [15]’s method. Our implementation uses a key dictionary mapping from each of the 24 keys (12 major and 12 minor) to a 12-dimensional binary vectors, where 1s indicate scale notes. Key detection is performed by computing dot products between these vectors and a song’s pitch histogram, with the highest-scoring key determining the transposition needed to normalize to C major or A minor. Further details can be found in the code.
B.2 Random Track Deletion
To encourage creativity in band arrangement and prevent the model from over-relying on direct copying notes from the input, we introduce a random track deletion mechanism during training. The intuition is that a well-trained model should be able to infer suitable instrumental content even if such content does not exist in the original music composition.
Concretely, given a content sequence with multiple instrument tracks, we randomly delete a subset of instruments before feeding the sequence to the model. The number of instruments to delete is sampled from a Poisson distribution with $\lambda=\max(\lfloor|\mathcal{I}|/4\rfloor,1)$, where $\mathcal{I}$ is the set of non-melodic instruments in the content sequence. The melodic track is estimated by the track with highest average pitch, and does not involves in the track deletion to ensure the melody’s content is preserved. Then, the deletion count is clipped to ensure at least one instrument remains. A corresponding number of instruments are uniformly sampled without replacement and all tokens associated with these instruments are removed from the input. Then the content sequence is calculated on this modified music sequence.
Importantly, the output target remains unchanged: it includes all instruments, including those that were deleted from the input. This setup requires the model to reconstruct the missing tracks based solely on the remaining musical context. By doing so, the model learns not only to replicate existing content, but also to infer what kind of notes would be appropriate for a given instrument in a specific musical context, even when such content is not explicitly present in the input.
B.3 Instrument Quantization
@@ B.4 Baseline Models
When comparing the effectiveness of REMI+ and the proposed REMI-z tokenization, the model structure used is the same as our arrangement model, but with a simpler training objective—the standard left-to-right next-token-prediction.
@@
For band arrangement, we use the Transformer-VAE model from [39] as our baseline, utilizing their official implementation without retraining. The model’s generation module was trained on Slakh2100 [19] (the same dataset as our fine-tuning) and its prior model leveraged the larger Lakh MIDI Dataset [22], which encompasses Slakh2100.
@@
Piano reductions can be categorized into two types: (1) piano accompaniments where the melody is delegated to a separate lead instrument, focusing solely on preserving harmony and texture (e.g., [38, 30]), which can be effectively handled using self-supervised methods; and (2) complete solo piano arrangements that additionally include the melody, requiring careful human arrangement and band-to-piano parallel datasets for supervised learning (e.g., [27, 26]). Due to the lack of open-source parallel datasets, we focus on the accompaniment arrangement task in this paper and use the term piano reduction interchangeably. However, our methodology is inherently flexible and can also handle the second type of reduction if parallel datasets become available. Additionally, exploring the potential for full solo piano reductions using unsupervised approaches remains a valuable direction for future research.
For comparison, we reimplemented the UNet baseline from [27] and trained it on Slakh2100 using our data preparation pipeline. Instead of adopting parallel data, we followed the same setting as our model, using the original composition as input and the piano track within the song as output. To ensure musical coverage, we selected only piano tracks that span more than 40% of the piece’s pitch range.
Since hand-specific annotations (i.e., left and right hand separation) are unavailable, the entire piano track is generated as a single sequence rather than as separate streams for each hand. We also omitted octave shifts during input preprocessing for two reasons. The first is that, unlike their setup where human-composed piano references may intentionally transpose notes by octaves, our output consistently corresponds to a subset of the input notes without such shifts. The second is empirical: introducing octave shifts during training significantly degraded UNet’s performance, reducing note-level F1 from 58.31 to 42.80 and diminishing perceptual quality.
@@
For drum arrangement, we employ Composer’s Assistant v2 [17] as our baseline, specifically using v2.1.0 from their official repository [^2] https://github.com/m-malandro/composers-assistant-REAPER. We use the model without retraining since its original training data (Lakh MIDI Dataset [22]) encompasses our fine-tuning dataset (Slakh2100).
B.5 Hyperparameter Settings
Table 7: Pre-train hyper-parameter setting.
@@ B.5 Hyperparameter Settings
For fine-tuning, we conducted a simple learning rate search over 1e-5, 5e-5, 1e-4, selecting the optimal value based on validation loss. This resulted in learning rates of 5e-5 for drum arrangement and 1e-4 for band arrangement and piano reduction. The batch sizes and context lengths were configured as follows: band arrangement used a batch size of 24 and context length of 768; piano reduction similarly adopted a batch size of 24 and context length of 768; drum arrangement employed a batch size of 8 and context length of 1536. Across all fine-tuning tasks, we used the AdamW optimizer with 0.01 weight decay, incorporating a linear learning rate scheduler with 500-step warmup. Training spanned 3 epochs for band, piano, and drum tasks, with early stopping patience of 2 epochs. The best checkpoints were selected based on validation loss.
@@ Appendix C Objective Metrics
In this section, we formally define the objective metrics used in the paper. All objective metrics are calculated at the segment level, and arithmetically averaged across the test set.
C.1 Arrangement Evaluation
Most of our objective metrics are based on note-level F1 scores calculated on piano roll of 16-th note quantization. Given the extreme sparsity of note events in the track-wise piano roll (e.g., only 0.12% non-zero elements in the Slakh2100 dataset under 16th-note quantization), F1-based metrics are more suitable than accuracy-based metrics for similarity evaluation. Let $X=\{x_{1},...,x_{n}\}$ and $Y=\{y_{1},...,y_{m}\}$ be two sequences of note events, where each note event $e$ is defined as a tuple $e=(t,p)$ with onset time $t$ and pitch $p$. A note event $x_{i}\in X$ is considered to match $y_{j}\in Y$ if and only if:
@@ C.1 Arrangement Evaluation
$\text{I-IoU}=\frac{|\mathcal{I}^{out}\cap\mathcal{I}^{tgt}|}{|\mathcal{I}^{out}\cup\mathcal{I}^{tgt}|}$ (9)
@@
$\text{V-WER}=\frac{S+D+I}{N}$ (12)
where $S$, $D$, and $I$ are the minimum number of substitutions, deletions, and insertions required to transform $V^{out}$ into $V^{tgt}$, and $N=|V^{tgt}|$ is the length of the $V^{tgt}$.
C.2 Tokenization Evaluation
Average tokens per bar ($\bar{T}_{\text{bar}}$) and average tokens per note ($\bar{T}_{\text{note}}$) are computed by first counting the total number of tokens and the number of bars or notes, respectively, across the entire dataset. Specifically, $\bar{T}_{\text{bar}}$ is obtained by dividing the total number of tokens by the number of bars, while $\bar{T}_{\text{note}}$ is obtained by dividing the total number of tokens by the number of notes.
Bar-level Shannon Entropy ($\bar{H}_{\text{bar}}$) is computed by measuring the Shannon entropy of each bar-level token sequence $H(X)$:
$H(X)=-\sum_{i=1}^{N}P(x_{i})\log_{2}P(x_{i}),$ (13)
where $X$ is the token distribution within a bar. Then we average it across all bars in the dataset to obtain $\bar{H}_{\text{bar}}$.
Note-level Perplexity ($\text{PPL}_{\text{note}}$) aggregates probabilities over all tokens representing a musical note (e.g., instrument, pitch, position, duration), and normalizes by the number of notes:
where $M$ is the number of notes in a bar.
Finally, token-level Perplexity ($\text{PPL}_{\text{token}}$) measures the standard autoregressive perplexity over all tokens. While commonly reported in language modeling, it serves as an auxiliary metric here, as it does not directly reflect the model’s ability to model note events, which are basic units of music.
@@ D.2 Human Evaluation Details
Survey Design. We conducted subjective evaluations using audio clips arranged by different models across three tasks: band arrangement, piano reduction, and drum arrangement. For band arrangement, we used out-of-domain test songs with novel compositions and instrument groups; for the other two tasks, test set songs were used to allow comparison with the original piano (Rule-O) or drum (ground truth) tracks. Each model was tasked with arranging the full song, and we selected a 15–30 second chorus phrase—the most representative segment—for evaluation. In total, we prepared 6 songs for band arrangement, 5 for piano reduction, and 5 for drum arrangement.
Each sample group includes the original music to be arranged and the anonymized outputs from all compared models (plus the ground truth for the drum task). Each sample is 8–16 bars long, rendered to audio in Cubase AI 13 using the default soundfont and the original MIDI’s BPM. We collected 56, 73, and 77 evaluation groups for band, piano, and drum tasks respectively. A total of 26 participants joined the study, 5 of whom work in the music industry. Among all evaluators, 73.1% have over 10 years of experience in music composition or performance. The mean time spent per session was approximately 30 minutes. Figure 6 shows the sample survey interface and instructions.
(a) An example of the piano reduction task.
(b) An example of the drum arrangement task.
Figure 6: Screenshots of survey pages and instructions of our online survey.
Evaluation Protocol. Participants first listened to the original music clip, followed by model outputs presented in random order. For each arranged clip, participants rated multiple aspects using a 5-point Likert scale, based on task-specific questions in the previous subsection. The evaluation groups were randomly distributed to participants to ensure unbiased feedback.
Band Arrangement Settings. To evaluate generalization, we designed three distinct instrumentation settings: (1) string trio (violin, viola, cello), (2) rock band (synth lead, clean electric guitar, distorted electric guitar, electric bass), and (3) jazz band (saxophone, violin, brass section, clean electric guitar, piano, string ensemble, electric bass). Each setting was applied to two different songs, covering six songs in total. The input content was drawn from both piano arrangements (3 songs) and band arrangements (3 songs), testing the model’s ability to adapt to varying content and target instrument groups.
@@ Appendix E Probing Analysis of Pre-Training Impact
Knowledge probing techniques are used to explore what a model has learned within its hidden representations [24]. A widely adopted method is linear probing, where the pre-trained model is frozen and a simple linear classifier is trained on top of the hidden representations to predict specific properties, thereby revealing the extent to which particular knowledge is encoded in the model [1]. This technique is particularly useful in our context for evaluating what the model has internalized during pre-training.
We have shown in the paper that the proposed models outperform the baseline models that do not undergo the pre-traning stage on the generation quality. In this section, we further analyze the reason. Specifically, we conduct probing experiments to assess whether pre-training enhances the acquisition of musical knowledge to facilitate understanding content conditions in fine-tuning.
We focus on two probing tasks: (1) classifying instrument types from content sequences that contain only position, pitch, and duration tokens, and (2) recognizing chord progression sequences from content sequences. To determine whether instrument or chord information is linearly accessible within the model’s sequence embeddings, we employ linear classifier probes. In these probing experiments, the average pooling of the Transformer’s output embeddings across all tokens in a sequence is fed into the classifier. The model parameters are kept frozen, and only the linear classifiers are trained. We compare the knowledge captured by different models: a randomly initialized Transformer, a model that has undergone pre-training only (PT only), a model trained on the band arrangement task without pre-training (FT only), and a model that has undergone both pre-training and fine-tuning (PT + FT).
E.1 Instrument Type Probing
Table 8: Instrument probing results.